2, 7, 26, 101, 400, ?
Answer by Edwin McCravy(20056)


2, 7, 26, 101, 400,

As a start, let's divide each term by the 
term before it: 

2nd term divided by 1st term:  7÷2 = 3.5
3rd term divided by 2nd term:  26÷7 = 3.7142... 
4th term divided by 3rd term:  101÷26 = 3.8846... 
5th term divided by 4th term:  400÷101 = 3.9603...

It looks like those answers are getting closer
and closer to 4.

So let's multiply all the terms by 4, and compare
what we get to the next term:

2×4 = 8, which is 1 more than the 2nd term 7
7×4 = 28, which is 2 more than the 3rd term 26
26×4 = 104, which is 3 more than the 4th term 101
101×4 = 404, which is 4 more than the 5th term 400

So we have found a pattern:.  Can you now figure out
the next term?  If not, tell me about it in the thank
you note form below and I'll get back to you by email.

Edwin